You own a $50,000 portfolio that is invested in a risk-free security and Stock A. The beta of Stock A is 2.20 and the portfolio beta is .95. What is the amount of the investment in Stock A?
Answer #1

Let investment in A=$x

Hence investment in risk free security=(50,000-x)

portfolio beta=respective beta*Respective weight

0.95=(x/50,000*2.2)+(50,000-x)/50,000*0[Beta of market=1;Beta of risk-free assets=0]

(0.95*50,000)=2.2x

x=47500/2.2

=$21590.91(Approx)=investment in Stock A.